OpenStudy (anonymous):

|7x-4|>= 3 help pls :)

sam (.sam.):

Do you have any ideas to nail this question?

OpenStudy (anonymous):

naaah:p

sam (.sam.):

You could square both sides and then solve for x \[|7x-4| \ge 3 \\ (7x-4)^2 \ge 9\] Try start here

OpenStudy (anonymous):

You have two inequalities to solve:\[ 7x-4\geq 3\quad\quad -(7x-4)\geq 3 \]

OpenStudy (anonymous):

thanx both :D